### v1

# Prioritized Trajectory Replay: A Replay Memory for Data-driven Reinforcement Learning

Abstract
In recent years, data-driven reinforcement learning (RL), also known as offline RL, have gained significant attention. However, the role of data sampling techniques in offline RL has been overlooked despite its potential to enhance online RL performance. Recent research suggests applying sampling techniques directly to state-transitions does not consistently improve performance in offline RL. Therefore, in this study, we propose a memory technique, (Prioritized) Trajectory Replay (TR/PTR), which extends the sampling perspective to trajectories for more comprehensive information extraction from limited data. TR enhances learning efficiency by backward sampling of trajectories that optimizes the use of subsequent state information. Building on TR, we build the weighted critic target to avoid sampling unseen actions in offline training, and Prioritized Trajectory Replay (PTR) that enables more efficient trajectory sampling, prioritized by various trajectory priority metrics. We demonstrate the benefits of integrating TR and PTR with existing offline RL algorithms on D4RL. In summary, our research emphasizes the significance of trajectory-based data sampling techniques in enhancing the efficiency and performance of offline RL algorithms.
Machine Learning, ICML
+jyliu@tju.edu.cn
1 Introduction
Reinforcement learning (RL) has made significant developments in recent years and has been widely applied in various scenarios (Ye et al., 2020; Kiran et al., 2022; Salvato et al., 2021). From a functional perspective, researchers usually study efficient RL algorithms from three aspects (Hessel et al., 2018): data collection, data sampling, and training algorithms. Data collection encompasses various methods for obtaining or generating diverse and comprehensive data, such as exploration (Hao et al., 2023) and data augmentation (Laskin et al., 2020). Data sampling refers to studying different sampling schemes based on existing data, to improve the learning efficiency (Schaul et al., 2016; Lee et al., 2019). Training algorithms are what most researchers focus on, aiming to maximize the optimization objectives through various techniques, including $Q$-value estimation (Mnih et al., 2015), representation (Sang et al., 2022), and more.
To overcome the high cost of interacting with the environment in real scenarios, offline RL has received wide attention as a way to learn good policy based on the fixed dataset. The study of offline RL can also be classified into the aforementioned three categories, but with slight differences. In offline RL, the literature mainly focuses on training algorithms through conservative estimation (Kumar et al., 2020; Bai et al., 2022), network structure (Chen et al., 2021), and other aspects. In terms of data collection, learning only from fixed dataset does not require exploration for new data, and research in data augmentation receives certain attention (Wang et al., 2022). Nevertheless, research on data sampling techniques under offline settings is still in its early stages and lacks a unified and comprehensive conclusion (Fu et al., 2022). Experience from online RL reveals that data sampling techniques can improve performance (Schaul et al., 2016). Therefore, this study primarily focuses on data sampling techniques to gain a more comprehensive understanding of their role in offline RL.
Figure 1: A motivating example on finite data. Left: The illustration of state transitions in three trajectories $\tau_{i}$ started at state $s_{0}$, with the reward $r$ for labeled state or 0 for others, and four different sampling techniques. Middle and right: Curves of the estimated maximum $Q$-value at state $s_{0}$ learned on these three trajectories. The solid line is the averaged value over 50 seeds, and the shaded area the standard variance. The oracle value, taking into account the discount factor, is slightly less than 8.
We first present an illustrative example in Figure 1 to analyze the effect of data sampling techniques on limited data. The results indicate that sampling state-transitions directly, whether being uniform (uni_state) or prioritized (prio_state), results in inferior learning of the $Q$-value at the initial state $s_{0}$. Even prioritized sampling (prio_state) offers little advantage over uniform sampling, which aligns with the conclusion of (Fu et al., 2022). Such sampling could potentially overlook critical states (such as $s_{1}$) that link other states to the initial state $s_{0}$, leading to ineffective propagation of subsequent reward signals to $s_{0}$. In contrast, when sampling along trajectories (uni_traj and prio_traj), the propagation of reward signals to $s_{0}$ is accelerated, converging faster to the optimal value. Such advantage is more prominent on sparse reward trajectories, where reward is given only at the last state. Compared to uniform trajectory sampling (uni_traj), prio_traj efficiently samples the trajectory with higher returns, further enhancing training efficiency. Our observation suggest that it is vital to consider the entire trajectory when determining the order of data sampling. Such trajectory perspective provides more informative insights into offline RL, rather than solely focusing on individual $(s,a)$ pairs.
To fully discover the potential of data sampling techniques from trajectory perspective for offline RL algorithms, we propose a plug-and-play memory, Prioritized Trajectory Replay (PTR), and conduct a detailed experimental analysis on these techniques, ultimately improving the performance of existing offline RL algorithms. We begin by implementing Trajectory Replay (TR), the basic memory to store and sample data in trajectories. Specifically, we sample and evaluate the states within a trajectory in a backward order. For instance, given two consecutive states, the latter one is sampled and evaluated first, followed by the former one. This allows each state to benefit from the estimated subsequent states, leading to improved convergence speed and performance especially on tasks with sparse rewards.
Based on TR, we continue to explore the potential of data sampling from two perspectives. From the algorithm updating perspective, we are inspired by EBU (Lee et al., 2019) and modify the target $Q$ computation method by balancing the original target $Q$-value and SARSA $Q$ target value (Sutton & Barto, 2018). The SARSA $Q$-value is calculated to limit the update process to use only trajectory data, avoiding the sampling of out-of-distribution (OOD) actions, thus alleviates the extrapolation error in offline learning. From the trajectory sampling perspective, we further propose Prioritized Trajectory Replay (PTR), considering the characteristics of trajectories (e.g., quality or uncertainty) and verifies different trajectory priority definitions. Through sufficient evaluation, we discover that prioritizing the sampling of trajectories with higher upper quartile mean or mean value of rewards, or trajectories with lower uncertainty, is effective in accelerating the learning of offline RL algorithms.
The main contribution of this paper is a concrete analysis of several trajectory-based data sampling techniques in offline RL, which are integrated into a unified memory structure - PTR. We emphasize that this work does not aim to propose a universally state-of-the-art (SOTA) algorithm. Rather, our study examines the advantage of basic trajectory sampling in TR, and the role of various trajectory priorities in PTR. We hope that our findings can provide useful insights towards developing powerful data sampling techniques in the future.
2 Related Work
This section provides an overview of fundamental data sampling methods and offline RL algorithms, which inspire our considerations of trajectory-based priorities in offline RL.
Data sampling in online RL
Data sampling studies how to efficiently sample data to improve training. In RL, this can be traced back to (Moore & Atkeson, 1993), which improves planning efficiency by selecting the next updated state based on priorities. For model-free RL, PER (Schaul et al., 2016) proposes to replay transitions with high temporal-difference (TD) error more frequently, which significantly improves the training efficiency of DQN (Mnih et al., 2015). Based on PER, several algorithms have been proposed, which optimize the error-based sampling (Pan et al., 2022; Li et al., 2021a), or propose other priority metrics (Oh et al., 2018; Fu et al., 2022), etc. These algorithms mostly focus on priority-based state-transition sampling. Other approaches study data sampling methods from the perspective of trajectory, and the most typical ones are EBU (Lee et al., 2019) and TER (Hong et al., 2022), which propose backward sampling from the end to the beginning of the trajectory. Backward sampling can make the update process more timely and accurately utilize the reward signals of the subsequent states to achieve accelerated learning.
Offline RL
In recent years, a large amount of research has focused on offline RL algorithms that learn promising policies from a fixed offline dataset. A key issue that offline RL algorithms need to address is how to avoid the impact of estimation errors (extrapolation errors) of unseen data in the dataset on policy learning, since collecting new data is impossible (Levine et al., 2020; Fujimoto et al., 2019). Common solutions can be divided into several categories: the first category imposes a policy constraint, which constrains the distribution of learned policies to be similar to the distribution of behavioral policies in the dataset (Kumar et al., 2019; Fujimoto & Gu, 2021); the second category applies support constraint, which require the action values of the learned policy to be restricted to the actions that appear in the dataset (Kostrikov et al., 2022; Xiao et al., 2023); the third category involves conservative $Q$-value (Kumar et al., 2020; Li et al., 2021b; An et al., 2021; Bai et al., 2022) or MDP model (Yu et al., 2021) estimation, which avoids selecting OOD actions during the policy update process by conservatively estimating unseen data.
In offline training, most research focuses on designing training algorithms, with less attention given to data sampling techniques. The literature (Fu et al., 2022) evaluates several data sampling techniques from a traditional state-transition perspective based on TD3+BC, but consistent conclusions have not been obtained. Considering the return of trajectories, ReD (Yue et al., 2022) uses return-based data re-balance for to improve the sampling probability of transitions on trajectories with higher returns. Determining the sampling order based on the trajectory return can be regarded as a special case of our study. Under the perspective of trajectory, we approach data sampling from two aspects: backward sampling on trajectories, and prioritized trajectory sampling based on trajectory quality or uncertainty.
3 Preliminary
3.1 Offline Reinforcement Learning
Offline RL refers to RL conducted on offline data. In offline RL, we have collected several trajectories, called the dataset $\mathcal{D}=\{\tau_{j}\}_{j=1}^{N}$. Each trajectory is a sequence of state-action-reward with length $l$: $\tau=\{(s_{i},a_{i},r_{i})\}_{i=0}^{l-1}$, where $s_{i}$ represents the state at time $i$, $a_{i}$ represents the action taken by the agent at that state, $r_{i}$ represents the immediate reward, and $s_{i+1}$ represents the next state. The optimization objective of offline reinforcement learning is usually to maximize the expected cumulative rewards through the dataset $\mathcal{D}$, without collecting new data.
3.2 Prioritized Experience Replay
Off-policy RL algorithms store historical experience in a replay memory. In traditional training, training data is uniformly sampled from the replay memory (Mnih et al., 2015). PER (Prioritized Experience Replay) (Schaul et al., 2016) prioritizes sampling data from the memory based on the TD-error of state-transitions, leading to significant improvement in DQN training efficiency. Specifically, the distribution $P(j)$ used for sampling transition $j$ is defined as follows:
$P(j)=$ $\frac{p_{j}^{\alpha}}{\sum_{k}p_{k}^{\alpha}},s.t.,p_{k}=|r_{k}+\gamma_{k}Q_{\mathrm{target}}(s_{k+1},$ (1)
$\arg\max_{a}Q(s_{k+1},a))-Q(s_{k},a_{k})|+\epsilon$
where $p_{j}$ is the priority of transition $j$, $\alpha$ determines how much prioritization is used, $\gamma$ is the discount factor, $Q(\cdot)$ and $Q_{\mathrm{target}}(\cdot)$ represent the $Q$-value and target $Q$-value for state-action pair $(s,a)$, and $\epsilon$ is a small constant that prevents priorities from being zero. In this way, PER can prioritize high-priority transitions based on the scale of TD-error, thereby accelerating the agent’s learning process.
4 Trajectory Replay
Figure 2: Overview of the process of data sampling based on Trajectory Replay.
In this section, we introduce Trajectory Replay (TR), a replay memory for storing offline data as complete trajectories and sampling batch data from a trajectory perspective. First, we describe the details of TR, including the technique of backward trajectory sampling used to sample from stored trajectories. We then introduce a weighted target computation based on TR, akin to EBU (Lee et al., 2019).
4.1 Trajectory Replay
To implement batch data sampling based on trajectory sequence, we maintain a basic replay memory storing the data in the form of trajectory, called Trajectory Replay, which differs from the traditional memory storing data as separate transitions. In EBU (Lee et al., 2019), it is suggested that sampling data in backward order on a trajectory is beneficial for online learning. Therefore, in this paper, we adopt backward sampling along this insight, as the default sampling order of TR. Such sampling based on TR is expected to offer certain advantages by making more timely and comprehensive use of information from successor state transitions.
Figure 2 illustrates the process of data sampling based on TR, in which $|N|$ trajectories in the offline dataset are stored. The available trajectories set stores all trajectories that have not been sampled. At the beginning, we sample $|\mathcal{B}|$ trajectories from it and add them to the sampled trajectories set. At each time step, the last state transition of $|\mathcal{B}|$ trajectories are moved to the data batch $\mathcal{B}$ for algorithm training. When all transitions of a trajectory in the sampled trajectories set have been added to $\mathcal{B}$, a new trajectory is required to be sampled from the available trajectories set to to fill up the sampled trajectories set. In such basic TR-based data sampling, the critic update process is not changed. Given the data batch $\mathcal{B}$, a typical optimization objective following deterministic actor-critic algorithm (Silver et al., 2014) is as follows:
$\mathcal{J}(\theta)=\mathbb{E}_{(s_{t},a_{t},r_{t},s_{t+1})\sim\mathcal{B}}\left[\ \frac{1}{2}\left(Q_{\theta}(s_{t},a_{t})-Q_{\mathrm{tg}}(s_{t},a_{t})\right)^{2}\ \right],$ (2)
$Q_{\mathrm{tg}}(s_{t},a_{t})=r_{t}+\gamma Q_{\bar{\theta}}(s_{t+1},\pi(s_{t+1})),$
where $\theta$ and $\bar{\theta}$ represent the parameters of critic network and target critic network, respectively, and $\pi(\cdot)$ represents the policy distribution.
[…]
4.2 Weighted Backward Update based on TR
A key challenge in offline RL is the extrapolation error arising from $Q$-value estimation. Although TR implements a basic trajectory-based sampling process, it does not take into account alleviating the extrapolation error, which can limit offline RL performance. Our research shows that a SARSA-like target can be introduced based on TR’s trajectory backward sampling process to completely avoid accessing OOD actions when calculating TD error during training. The critic target in Eq. 2 is now as follows:
$Q_{\mathrm{target}}(s_{t},a_{t})$ $=r_{t}+\gamma Q_{\mathrm{target}}(s_{t+1},a_{t+1}),$ (3)
where $Q_{\mathrm{target}}(s_{t+1},a_{t+1})$ is the target $Q$-value, calculated in the next time step in the same trajectory. Optimizing Eq. 3 can be considered as an implicit support constraint as OOD actions are avoided.
Based on it, we modify the computation of the target $Q$ by balancing the original target in Eq. 2 and the vanilla SARSA target, to introduce the implicit support constraint in Eq. 3. We obtain the following weighted form of target estimation:
$Q_{\mathrm{target}}(s_{t},a_{t})=r_{t}+\gamma$ $[(1-\beta)Q_{\mathrm{target}}(s_{t+1},a_{t+1})$ (4)
$+\beta Q_{\bar{\theta}}(s_{t+1},\pi(s_{t+1}))],$
where $\beta$ is the hyper-parameter that controls how much the original target should be used.
5 Prioritized Trajectory Replay
Building upon Trajectory Replay, we have developed a weighted target to improve learning efficiency. In this section, we shift our focus from the updating process to enhancing the trajectory sampling process depicted in Figure  2. To enhance the performance through more properly sampling trajectories, we introduce several novel trajectory priority metrics to prioritize trajectories during sampling. These metrics primarily consider two key factors: the quality of the trajectory return and the degree of trajectory uncertainty.
Instead of uniformly sampling from the available trajectories, we introduce a probabilistic sampling approach that incorporates trajectory priority metrics for trajectory $\tau_{j}$, denoted as $\mathrm{pri}(\tau_{j})$, to prioritize trajectories. To avoid potential biases towards high priority trajectories, we use the ranking order, denoted as $\mathrm{rank}(\mathrm{pri}(\tau_{j}))$, as a measure to determine the sampling probability $p_{\tau_{j}}$, rather than relying solely on the absolute value of $\mathrm{pri}(\tau_{j})$. Assigning appropriate priority to each trajectory through its rank ensures that all available trajectories are able to be utilized effectively. To this end, we define the prioritized sampling probability distribution $P(\tau_{j})$ for all available trajectories as follows:
$P(\tau_{j})$ $=\frac{p_{\tau_{j}}^{\alpha}}{\sum_{k}p_{\tau_{k}}^{\alpha}},\ \text{s.t.},\ p_{\tau_{j}}=\frac{1}{\mathrm{rank}(\mathrm{pri}(\tau_{j}))}.$ (5)
5.1 Priority based on Trajectory Quality
Higher quality trajectories have been found to be more beneficial for offline RL (Fujimoto et al., 2019). To this end, we propose utilizing trajectory quality as a priority metric during the data sampling process to prioritize high-quality trajectories. Recognizing that using return as a metric may unfairly prioritize longer trajectories, we introduce alternative measures such as the mean of rewards on the trajectory, including upper quartile mean (UQM) and upper half mean (UHM). In this regard, we define six distinct metrics to assess trajectory quality, as follows:
  • Return. The undiscounted sum rewards of the trajectory.
  • Avg. reward. The averaged reward of the trajectory.
  • UQM reward. The averaged reward of the top 25% state transitions in the trajectory.
  • UHM reward. The averaged reward of the top 50% state transitions in the trajectory.
  • Min reward. The minimum reward of the trajectory.
  • Max reward. The maximum reward of the trajectory.
Trajectory priority $\mathrm{pri}(\tau_{j})$ is established based on these metrics to prioritize the sampling of high-quality trajectories. In particular, Return and Avg. reward reflect the overall performance of the trajectory, while UQM reward, UHM reward, and Max reward reflect the quality of the best transitions in the trajectory. on the other hand, Min reward prioritizes trajectories with higher minimum reward values, indicating a more stringent criterion for high quality.
5.2 Priority based on Trajectory Uncertainty
Within the context of offline RL, uncertainty serves as a pivotal metric by quantifying the level of disagreement present among $Q$ estimates. Such disagreement indicates a lack of confidence in the estimated $Q$-value for a given state-action pair. Thus, a higher degree of uncertainty indicates greater unreliability in the knowledge of the state-action pair, making it unwise and infeasible to use it for policy execution. Taking a trajectory perspective, we define trajectory priority based on uncertainty, such that trajectories exhibiting lower degrees of uncertainty are assigned greater priority during sampling. Specifically, we establish trajectory priority metrics based on the mean, upper quartile mean (UQM), and lower quartile mean (LQM) of all state-action pairs’ uncertainty values present within the trajectory, as follows:
  • Lower mean unc. The reciprocal of the average uncertainty of all state-action pairs.
  • Lower LQM unc. The reciprocal of the average uncertainty of bottom 25% state-action pairs.
  • Lower UQM unc. The reciprocal of the average uncertainty of top 25% state-action pairs.
Alternatively, we can assign higher priority to trajectories with higher levels of uncertainty. We can obtain the following trajectory priority metrics by taking the reciprocals of the aforementioned values: Higher mean unc., Higher LQM unc., and Higher UQM unc., respectively.
By incorporating trajectory priority, we upgrade TR to Prioritized Trajectory Replay (PTR), achieving probabilistic sampling of trajectories. We anticipate that prioritizing trajectories with higher quality would be more suitable for sparse reward tasks, where the reward signals exhibit significant variations on these trajectories and thus require an emphasis on trajectories with higher value for learning during probabilistic sampling. Also, we expect that prioritizing trajectories with lower uncertainty can be effective in dense reward tasks, which can aid in accurately estimating the level of uncertainty. Conversely, prioritizing trajectories with higher uncertainty is not expected to work well, as errors from data with high uncertainty can accumulate severely for finite data and hinder learning. Section 6 provides a comprehensive and detailed empirical evaluation of these distinct trajectory priority metrics.
6 Experiment
To reveal the consistency between our analysis and the performance of TR/PTR, and demonstrate the advantage over other advanced methods, we conduct experiments to address the following questions:
  • RQ1 (Trajectory replay): Can basic sampling of TR bring an improvement to offline RL algorithms by efficiently using information from successor state transitions?
  • RQ2 (Weighted target based on TR): Can modifying the critic target to SARSA-style target or weighted target bring further improvement to the performance, and how do they demonstrate?
  • RQ3 (Prioritized trajectory replay): Without changing the critic target, can the prioritized trajectory sampling bring consistent performance improvement, and what are the characteristics of different trajectory priority metrics?
6.1 Baseline Algorithms and Implementation Details
Benchmark
In order to comprehensively analyze the advantages and disadvantages of various proposed data sampling mechanisms from a trajectory perspective, we conduct empirical evaluations on the D4RL benchmark (Fu et al., 2020), focusing primarily on the Mujoco -v2 datasets (dense reward), Adroit -v1 datasets (sparse reward), and AntMaze -v0 datasets (sparse reward).
Baselines
Baseline algorithms include (i) TD3+BC, which uses regularization terms of behavior cloning to constrain the learning policy, (ii) IQL, which uses expectile regression to focus on in-sample actions and avoids querying the values of unseen actions, and (iii) EDAC, which implements conservative estimation of critic based on ensemble $Q$ networks. The most critical hyper-parameters in TD3+BC is $\alpha$ used to control the weights of RL and imitation learning. On Mujoco tasks, we set $\alpha$ to 2.5 and the reproduced results are similar to those reported in the original paper. On the Antmaze and Adroit datasets, we conduct a hyperparameter search in the ranges 0.0001, 0.05, 0.25, 2.5, 25, 36, 50, 100, resulting in better results than previously reported. The most important hyper-parameters of EDAC algorithm are the $n$ used to control the number of ensembles, and $\eta$, the weight of the ensemble gradient diversity term. To make it easier to verify the advantage, we set $n$ to 3 for mujoco-medium-replay-v2 datasets and 10 for others. We use the default hyper-parameters for IQL, as recorded in the original paper. All hyper-parameters mentioned above are listed in the Appendix.
Implementation details
To ensure code conciseness, readability, and a fair and identical experimental evaluation across algorithms, we reproduce the baseline algorithms based on the CORL repository (Tarasov et al., 2022) and implement various sampling techniques. Unless otherwise specified, all results reported in the following for the baseline algorithms are the results we have reproduced ourselves. For our methods, various data sampling techniques are implemented based on a plug-and-play memory module, TR/PTR. The only hyper-parameter $\beta$ is used to construct the weighted critic target, and we conduct grid searches mainly using values from 0.05, 0.25, 0.5, 0.75, 0.95, 0.98 in the experiments. In basic TR and trajectory-based priority sampling, no hyper-parameters needs to be fine-tuned. The supplementary materials include our code.
6.2 Evaluation of Trajectory Replay
Table 1: Normalized average returns of TR and baselines on Gym Mujoco, Antmaze, Adroit tasks in D4RL. The averaged performance and standard deviation of 3 runs are reported.
Figure 3: Comparison of the performance of PTR compared to TD3+BC under 10 different trajectory priority metrics. In the figure we restrict the performance difference to a maximum of 20.
Task Name TD3+BC TD3+BC
EDAC EDAC
IQL IQL
Mujoco
659.81
630.54
617.71
546.44
691.41
669.99
Antmaze
98.36
223.67
329.90
356.47
Adroit
540.41 545.95
2.37
23.41
545.70 543.52
We first evaluate our proposed TR by implementing the most basic trajectory perspective-based data sampling technique with trajectory backward sampling, which serves as the basis for follow-up on critic target and priority trajectory sampling. We present the reproduced results of the baseline algorithm and the performance improvements achieved by combining TR with these algorithms in Table 6.2, and full results of all datasets in Appendix.
The results suggest that backward sampling based on TR is particularly effective in tasks with sparse rewards. Consistent performance improvements of TR are shown in the Antmaze and Adroit datasets, with notable score increases achieved by TD3+BC, EDAC, and IQL on Antmaze of over 100, 20, and 30 points respectively. These findings support the efficacy of trajectory backward sampling as an approach for efficiently utilizing reward information from subsequent state transitions, which is consistent with previous research on online scenarios.
However, for dense reward tasks, the benefits of TR are limited. On datasets walker2d-medium-replay-v2 and hopper-medium-replay-v2, there is even a noticeable decrease in performance. This highlights the urgent need for deeper investigate of data sampling techniques based on TR to improve offline RL algorithms’ performance.
On dense-reward datasets, prioritizing trajectories with lower uncertainty is reasonable to achieve stable and outstanding performance, as demonstrated by the last three subfigures in Figure 6.2. These metrics gain better or equivalent performance on Mujoco datasets. This highlights the ability of uncertainty to capture the characteristics of dense reward trajectories. The metric Lower UQM unc., which seeks trajectories with lower overall uncertainty, is the best performing metric in these cases.
Additionally, for dense reward tasks, when there are significant differences in rewards between trajectories, such as mujoco-medium-expert dataset, prioritized sampling based on trajectory quality is also a good option, which can prioritize the use of higher-quality data and achieve better results. However, this approach may harm performance on the mujoco-medium-replay dataset. On the other hand, estimating uncertainty of data with sparse reward signals can be challenging, and thus, the performance of uncertainty-based metrics is limited on sparse-reward datasets.
Table 4: The comparison of various priority metrics
The dataset property
TR
Uniform
PTR
Quality
PTR
Uncertainty
Sparse reward trajectories
Dense reward trajectories
Many but not all high quality trajectories
In summary, as shown in Table 4, for tasks with sparse rewards, the basic trajectory sampling schema in TR can already show advantages, and it is also advisable to prioritize trajectories based on their quality, especially prioritizing based on UQM or Avg. reward, to gain better overall performance. Besides, for tasks with dense rewards, prioritizing trajectories of lower uncertainty is recommended, with Lower UQM unc. being the best metric for greater benefits. If there are ample high-quality data in the dataset, prioritizing based on trajectory quality is also worth considering.
Figure 3: Comparison of the performance of PTR compared to TD3+BC under 10 different trajectory priority metrics. In the figure we restrict the performance difference to a maximum of 20.